The letters of the word "SIXTEEN" are randomly arranged. What is the probability that the two E's are not next to each other?
Answer by robertb(5830)

# of arrangements the two E's are together = 6!1! = 720
# of arrangements of the letters of the word SIXTEEN = 7!/2! = 2520
probability that the two E's are together = 720/2520 = 2/7.
Therefore probability that the two E's are not next to each other = 1-2/7 = 5/7.